OpenStudy (johnw):

Determine a primitive function to f(x) = e^-x + 1/√1-x^2 + cos(πx)

OpenStudy (eliesaab):

\[ f(x)= e^{-x}+ \frac{1}{\sqrt{1-x^2}}+\cos (\pi x) \]

OpenStudy (eliesaab):

Is your function like the above? Take the primitive of each one of them and add them up

OpenStudy (eliesaab):

The first and the last one are easy. The second one think about a inverse trig function

OpenStudy (eliesaab):

It seems that you are not here. Anyway, you have enough hints to finish your problem

OpenStudy (johnw):

yes, it´s the same

OpenStudy (johnw):

I´m here, but I didn´t get any response from anybody for like an hour so I did something else :)

OpenStudy (eliesaab):

What is a primitive of \( e^{-x}\)?

OpenStudy (eliesaab):

Remember that the derivative of \( e^{ a x}\) is \( a e^{ax}\) where is a constant.

OpenStudy (johnw):

so the primitive of e^-x is -e^-x? Or is it + a constant?

OpenStudy (johnw):

Time for bed. Here in Sweden it´s half past midnight :) I´ll leave the question open....